# Classifying representations of finite classical groups of Lie type of dimension up to $\ell^{4}$

### v1

@@ Abstract.
In this article we classify the irreducible representations of finite classical groups $G$ of Lie type of rank $\ell$ with dimension at most a constant proportional to $\ell^{4}$ according to whether $G$ is of type $A_{\ell}$ or not. The dimensions of those representations are explicitly computed by hand. We conclude the work by addressing certain cases of the inverse Galois problem via the foregoing classification.
@@ 1. Introduction
It is well-known the multiples bridges among representation theory with other theories in mathematics. We are specially concerned about the irreducible representations of finite classical groups of Lie type and its connections with the large image conjecture for automorphic Galois representations and the inverse Galois problem. More precisely, let $G_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$ and $\mathbb{A}_{\mathbb{Q}}$ the ring of adeles of $\mathbb{Q}$ and $\pi$ be a cuspidal algebraic regular automorphic representations of $\operatorname{GL}_{n}(\mathbb{A}_{\mathbb{Q}})$. According to global Langlands correspondence [5], for each prime $p$ there should exist a $p$-adic Galois representation $\rho_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{GL}_{n}(\overline{\mathbb{Q}}_{p})$ associated to $\pi$. A folklore conjecture (usually known as the large image conjecture) predicts that the image of the proyectivization $\overline{\rho}^{proj}_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{PGL}_{n}(\overline{\mathbb{F}}_{p})$ of the residual representation $\overline{\rho}_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{GL}_{n}(\overline{\mathbb{F}}_{p})$ of $\rho_{\pi,p}$ should be as large as possible for almost all primes $p$ (i.e., all but finitely many), except when an automorphic reason impede it. For instance, when $\pi$ is polarizable [2, §2.1], this means that the image of $\overline{\rho}^{proj}_{\pi,p}$ should contain either $\operatorname{PSp}_{n}(\mathbb{F}_{p^{s}})$ or $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}})$ (for some positive integer $s$) for almost all $p$. This conjecture is, in general, known in very few cases ($n\leq 6$) and the proofs in all known cases when $\pi$ is polarizable are done by considering all possible images of $\overline{\rho}^{proj}_{\pi,p}$ given by the classification of the maximal subgroups of $\operatorname{PSp}_{n}(\mathbb{F}_{p^{s}})$ or $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}})$see [15] for more details. As we will see at the end of this paper, this conjecture has important consequences on the Inverse Galois problem for finite groups of Lie type.
@@
We would like to point out that we are assuming $G$ of simply connected type, which ensures a correspondence between representations of the group $G$ and representations of its Lie algebra $\mathfrak{g}$ for almost all $p$. Therefore, through this paper, we will consider $p$ sufficiently large in order to account with such a correspondence. The main result of this paper splits into three theorems which tackle the classification of irreducible representations of $G$ of dimension less or equal to $\ell^{4}/16$ or $\ell^{4}$ according to whether the group is of type $A_{\ell}$ or not. Thanks to the work of Chevalley, we know that each irreducible $G(\overline{\mathbb{F}}_{p})$-module $V$ is determined (up to isomorphism) by its highest weight and for each dominant weight $\lambda$ there is an irreducible $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ with highest weight $\lambda$. Thereby, classifying those irreducible representations is equivalent to determine all dominant weights $\lambda$ such that the dimension of the irreducible $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ of highest weight $\lambda$ satisfies the foregoing condition-we will call them admissible weights. Determining this class of weights will imply to compute diverse dimensions and consequently to handle a considerable number of multiplicities. We then will present certain examples with these calculations first to illustrate admissible and not admissible weights. Nevertheless, one of the key motivations to give these examples is to reminisce about the computing of multiplicities and its recursive nature since we will have to compute longer multiplicities and dimensions along the work. In addition, a second good motivation is that those examples are placed in such a manner to contribute in the proofs of our main results, helping to be more readable such demonstrations.
The classification of admissible weights splits into three stages according to $G$ is of type $A_{\ell}$, $B_{\ell}$ or of type $C_{\ell}$ and $D_{\ell}$. Our strategy to this end will be to analyze the behavior of diverse multiplicities of weights $\mu\preceq\lambda$ to study the dimension of $V(\lambda)$ whereby we obtain a good list of admissible weights in order to show that this list exhaust the admissible ones. We therefore commence by applying successfully the mentioned strategy to $G$ of type $A_{\ell}$ as starting case. Later on, we tackle the case when $G$ is of type $B_{\ell}$, where we may point out that determining the admissible weights certainly presents additional challenges. Indeed, computing the required multiplicities to study admissibility of weights implies much more efforts than the case $A_{\ell}$. In particular, these calculations turn out to be very laborious with extensive computations to use the Freudenthal formula to determine multiplicities at hand. We are able to carry out all those computations and find a good list of admissible weights. Once we have ended that, obtaining good lists for types $C_{\ell}$, $D_{\ell}$ will follow more natural.
The second stage of the strategy also will be a subtle task in the sense that from an arbitrary admissible weight we will have to prove this one can be reduced to a one of our good list, i.e., the list exhausts the admissible ones which gives rise to the desired classification. We here continue case-by-case tackling first the type $A_{\ell}$ and subsequently the case of type $B_{\ell}$. This work takes a significant time and whence the cases $C_{\ell}$ and $D_{\ell}$ will be deduced directly. Hence we only outline the proof by refining the arguments according to the case and by proving the admissibility of the new weight that appears when $G$ is of type $C_{\ell}$ and $D_{\ell}$. At this point, we would like to observe that the required multiplicities in these cases are listed in two tables of the Appendix, where we bring all information to compute them, thus avoiding repetitive arguments and calculations.
We conclude the Introduction by giving the structure of the paper.
This article begins with a brief background on representation theory of finite groups of Lie type in §2. Here we introduce the root systems and fix the orders that we will be used through the work. Also we recall the Freudenthal multiplicity formula and the expression to compute dimension of the $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ with highest weight $\lambda$. In §3 we introduce the concept of admissible weight $\lambda$ when $G$ is of type $A_{\ell}$ and obtain a good list -Proposition 3.6- which is the first part toward the classification of admissible weights of $G$. Later on, §4 is dedicated completely to Theorem 4.1, where the classification of the admissible weights is obtained when $G$ is of type $A_{\ell}$. In §5 we tackle our aim for groups $G$ of type $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$. In the first case we prove a series of lemmas on required multiplicities to obtain a list of admissible weights -Proposition 5.6. Thus, we may show such a classification for $G$ of type $B_{\ell}$ in Theorem 5.9. While Theorem 5.10 determines all the admissible weights for the remainder two types by appealing to Appendix -Tables 1 and 2- for the required multiplicities. We present the last two results of the article in §6 by providing applications of our classification to large image conjecture for automorphic Galois representation in Theorem 6.2 and on the inverse Galois problem in Corollary 6.3. We conclude our work with an appendix §7 where we give the tables with the all required multiplicities when $G$ is of type $C_{\ell}$ or $D_{\ell}$.
@@ 2. Preliminaries on representations of finite simple groups of Lie type
In this section we will recall some results about algebraic groups and their defining characteristic representations that we will need in this paper.
@@ 2.1. Root systems
Let $G$ be a connected reductive simple algebraic group over $\overline{\mathbb{F}}_{p}$ of simply connected type and rank $\ell$. Let $T$ be a maximal torus of $G$, $X\cong\mathbb{Z}^{\ell}$ its character group and $Y\cong\mathbb{Z}^{\ell}$ its co-character group. Let $\{\alpha_{1},\ldots,\alpha_{\ell}\}\subset X$ be a set of simple roots for $G$ with respect to $T$ and $\Delta$ its associated root system. Let $\alpha_{i}^{\vee}\in Y$ the coroot corresponding to $\alpha_{i}$ for each $1\leq i\leq\ell$. Viewing $X\otimes\mathbb{R}$ as Euclidean space we can define the fundamental weights $\lambda_{1}\ldots,\lambda_{\ell}\in X\otimes\mathbb{R}$ as the dual basis of $\alpha_{1}^{\vee},\ldots,\alpha_{\ell}^{\vee}$.
For our purposes in accordance with [11] we shall use the opposite order from the one given in [3, 8]. Thus, the Dynkin diagrams for classical groups are given as follows:
We also recall that each $\alpha\in\Delta$ determines a reflection $\sigma_{\alpha}$ by means of
$\sigma_{\alpha}(\beta)=(\alpha,\beta),\quad\quad\beta\in E,$
where $(E,(-,-))$ is the underlying inner product real vector space of finite dimensional associated to $\Delta$. The Weyl group $\mathcal{W}$ of $\Delta$ is the subgroup of the orthogonal group on $E$ generated by the reflections $\sigma_{\alpha_{i}}$. Thus one observes that this group acts on $\Delta$ by permutation. In fact, $\mathcal{W}$ is generated by the reflections $\sigma_{\alpha_{i}}$ with $1\leq i\leq\ell$ by appealing to [9, Prop. 2.62].
Furthermore, one obtains a partial ordering in $X$ by declaring that a weight $\lambda$ is smaller than a weight $\lambda^{\prime}$ (denoted by $\lambda\prec\lambda^{\prime}$) if and only if $\lambda^{\prime}-\lambda$ is a non-negative linear combination of simple roots. We shall also use $\preceq$ in an obvious way. In addition, a weight $\lambda\in X$ is dominant if it is a non-negative linear combination of the fundamental weights. The group $\mathcal{W}$ acts on the set of the dominant weights by means of its action on the simple roots $\alpha_{1},\cdots,\alpha_{\ell}$. Given a dominant weight $\lambda$, we will denote by $\operatorname{Orb}_{\mathcal{W}}(\lambda)$ and $\rm{Stab}_{\mathcal{W}}(\lambda)$ the orbit and the stabilizer of $\lambda$ under the $\mathcal{W}$-action, respectively.
It is worth recalling that the stabilizer subgroup $\rm{Stab}_{\mathcal{W}}(\lambda)$ of a dominant weight $\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}$ is the parabolic subgroup of $G$ generated by the reflections along the simple roots $\alpha_{i}$ such that $a_{i}=0$.
@@ 2.2. Representations
Let $V$ be a finite-dimensional $G(\overline{\mathbb{F}}_{p})$-module. Considering this as $T$-module, we have a decomposition $V=\bigoplus_{\lambda\in X}V_{\lambda}$ into weight spaces $V_{\lambda}=\{v\in V:vt=\lambda(t)v\mbox{ for all }t\in T\}$. Then, the set of weights $\lambda\in X$ with $V_{\lambda}\neq\{0\}$ is called the set of weights of $V$. This set is a union of $\mathcal{W}$-orbits.
As we pointed out in the Introduction, from a result of Chevalley (see Section 31.3 of [8]) we have that, if $V$ is irreducible, then the set of weights of $V$ contains a unique element $\lambda$ such that for all weight $\lambda^{\prime}$ of $V$ we have $\lambda^{\prime}\prec\lambda$. This $\lambda$ is called the highest weight of $V$. This is dominant and $\dim(V_{\lambda})=1$. Moreover, we have that each irreducible $G(\overline{\mathbb{F}}_{p})$-module $V$ is determined (up to isomorphism) by its highest weight and for each dominant weight $\lambda\in X$ there is an irreducible $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ with highest weight $\lambda$.
For example, when $G$ is of $A_{\ell}$ type, the standard representation $V=V(\lambda)$ has highest weight $\lambda=\lambda_{\ell}$ and dimension $\ell+1$. Moreover, the anti-symmetric tensor product $\bigwedge^{k}(V)=V(\lambda)$ is irreducible with highest weight $\lambda=\lambda_{\ell-k+1}+\cdots+\lambda_{\ell}$ and dimension $\binom{\ell+1}{k}$. More generally, the irreducible representation $V(\lambda)$, with highest weight $\lambda=a_{1}\lambda_{1}+(a_{1}+a_{2})\lambda_{2}+\cdots+(a_{1}+\cdots+a_{\ell})\lambda_{\ell}$, is contained in
$\operatorname{Sym}^{a_{1}}(\Lambda^{\ell}V)\otimes\cdots\otimes\operatorname{Sym}^{a_{\ell-1}}(\Lambda^{2}V)\otimes\operatorname{Sym}^{a_{\ell}}(V)$
by Weyl’s Construction [7, Lecture 15]. In particular, taking $a_{\ell}\neq 0$ and $a_{i}=0$, $i=1,\cdots,\ell-1$, we have that $V(a_{\ell}\lambda_{\ell})$ is contained in the $a_{\ell}$-symmetric power $\operatorname{Sym}^{a_{\ell}}(V)$ of $V$, which has dimension $\binom{\ell+a_{\ell}}{a_{\ell}}$. In fact, we will prove that $\operatorname{Sym}^{a_{\ell}}(V)$ is irreducible by computing the dimension of $V(a_{\ell}\lambda_{\ell})$. In general, when $G$ is of type $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$, we can only guarantee that $V(k\lambda_{\ell})\subseteq\operatorname{Sym}^{k}(V)$ and $V(\lambda_{\ell-k+1}+\cdots+\lambda_{\ell})\subseteq\Lambda^{k}(V)$. Through this paper, we also compute the exact dimensions of these irreducible representations $V(\lambda)$ for $k=2,3,4$ by computing multiplicities as follows.
2.3. Multiplicity and dimension
First, we introduce a special weight which appears in computing of multiplicities as we will see here below. Specifically, the strongly dominant weight $\delta$ is defined by

### v3

@@ Abstract.
Let $G$ be a finite classical group of Lie type of rank $\ell$, defined over a field of characteristic $p>2$. In this work, we classify the irreducible representations of $G$ whose dimensions are bounded by a constant proportional to $\ell$, and splits into two cases according to $G$ is of type $A_{\ell}$ or not. Furthermore, we discuss explicit formulas for computing the dimensions of such representations. The motivation for this work arises, in part, from a desire to obtain new results on two classical problems concerning Galois representations: the large image conjecture for automorphic Galois representations and the inverse Galois problem. We conclude the paper by giving some remarks on potential implications in these addresses.
Keywords: Finite classical groups of Lie type, computing of multiplicities, dimensions of weight spaces, image of Galois representations.
@@ 1. Introduction
In this paper we are particularly interested in the representation theory of the finite classical groups $G$ of Lie type of rank $\ell$. It is well known that every irreducible representation $V$ of a simple connected, simple algebraic group over a field $K$ is in correspondence with a suitable dominant weight $\lambda$, called highest weight, of a root system associated to maximal torus $T$ of $G$. In particular, for any dominant weight $\lambda$, there exists an irreducible finite-dimensional $G(K)$-module $L(\lambda)$ of highest weight $\lambda$. This module is not fully understood for an arbitrary weight $\lambda$. For instance, differences in the dimension of $L(\lambda)$ are expected according to the characteristic of $K$: in characteristic zero, the dimension of $L(\lambda)$ is given by the Weyl’s formula, whereas in positive characteristic $p$, a closed formula for the dimension of $L(\lambda)$ is still far to be known for an arbitrary weight $\lambda$. In this case, the multiplicities involved in computing $\dim L(\lambda)$ depend on $p$, and determining this dependence is a challenging task. In this regard, progress have been established case-by-case according to $G$ and very special weights $\lambda$ (see [7, 8, 9, 13, 16, 23, 24, 32]-for instance). With this in mind, we aim to classify irreducible modules whose dimension is bounded by a given constant. More precisely, we will say that $\lambda$ is an admissible weight if
@@
A complete classification of all admissible weights for arbitrary $n$ turns out to be a major challenge, which makes a case-by-case approach both natural and plausible. This way was already started by Liebeck in [23], where all the admissible weights for $n=2$ were determined. Later, Lübeck [24] extended those results in the case of $n=3$, providing, in addition, a series of tables listing representations whose dimension is less than or equal to a bound $M$, depending on the type. An extension of such a classification for bounds $\binom{\ell+1}{4}$ or $16\binom{\ell}{4}$ according to the type is given by Martínez [25].
The main result of this paper split into three theorems, in which we classify all admissible weights when $n=4$ in Eq 1.1. More precisely, in §3, we state our first result -Theorem 3.4- concerning the classification of all the admissible weights when $G$ is of type $A_{\ell}$. The proof begins with an arbitrary dominant weight $\lambda$ and consists on determining all the possibilities of its coefficients via a series of reductions until to left reduced to a full list of admissible weights. Later, in §4 we provide the classification of admissible weights for groups $G$ of type $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$. Here, our method implies to obtain a series of bounds for the multiplicities for the weights involved in the classification. In this vein, for $B_{\ell}$, we have to prove a series of lemmas on required multiplicities in order to obtain a list of admissible weights -achieved in Proposition 4.4. We then deduce, in Theorem 4.6, that such list exhausts all the admissible weights by means of successive reductions. The section concludes with Theorem 4.7, where all admissible weights for the remaining two types are classified by appealing to Appendix -Tables 1 and 2- for the required multiplicities.
As noted above, determining the exact dimension of $L(\lambda)$ in positive characteristic is significantly more complicated than in characteristic zero, since a closed formula for this dimension is unknown. Here, computing the multiplicity $m_{\lambda}(\mu)$, for any dominant weight $\mu\preceq\lambda$, requires constructing a $d\times d$-matrix $C^{(\lambda,\mu)}$ and then the multiplicity of $\mu$ is the rank of the matrix obtained from $C^{(\lambda,\mu)}$ by reduction modulo $p$, where $d$ depends on the Kostant partition function. In particular, determining the precise dependence of the multiplicities involved in $\dim L(\lambda)$ upon reduction modulo $p$ relies heavily on the nature of $\lambda$, and demands substantial effort. In $\S 5$, we compile and supplement the dimension formulas of $L(\lambda)$ for all our admissible weights $\lambda$ when $p$ is sufficiently large.
Now, let $G_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$, and let $\pi$ be a cuspidal, algebraic regular, automorphic representations of $\operatorname{GL}_{n}(\mathbb{A}_{\mathbb{Q}})$, where $\mathbb{A}_{\mathbb{Q}}$ denotes the ring of adeles of $\mathbb{Q}$. According to global Langlands correspondence [11], for each prime $p$ there should exists a $p$-adic Galois representation $\rho_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{GL}_{n}(\overline{\mathbb{Q}}_{p})$ associated to $\pi$. A folklore conjecture predicts that the image of the projectivization $\overline{\rho}^{proj}_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{PGL}_{n}(\overline{\mathbb{F}}_{p})$ of the residual representation $\overline{\rho}_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{GL}_{n}(\overline{\mathbb{F}}_{p})$ of $\rho_{\pi,p}$ should be as large as possible for almost all primes $p$ (i.e., all but finitely many), except when an automorphic reason impedes it. This statement is often referred to as the (residual) large image conjecture and it has interesting consequences on the Inverse Galois problem for finite groups of Lie type [1, 2, 31, 36, 37, 38]. It is worth noting that in the few known cases ($n\leq 6$) of this conjecture for $\pi$ polarizable, its proof is done by considering all possible images of $\overline{\rho}^{proj}_{\pi,p}$ given by the classification of the maximal subgroups of $\operatorname{PSp}_{n}(\mathbb{F}_{p^{s}})$ or $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}})$.
In the eighties, Aschbacher [3] classified the maximal subgroups of almost all of the finite almost simple classical groups, dividing them into two classes. One of these consists of absolutely irreducible subgroups that arise (at least for sufficiently large $p$) as representations of finite groups of Lie type. However, this classification is not fine enough to contribute directly to the large image conjecture. The search for a more refined classification is precisely one of the motivations behind this work. We conclude this paper by showing a weak version of the large image conjecture, and obtain new weak results on the inverse Galois problem as illustrated in Theorem 5.3 and Corollary 5.4.
@@ 2. Preliminaries on representations of finite simple groups of Lie type
In this section, we recall some results about algebraic groups and their defining characteristic representations that will be needed in this paper.
@@ 2.1. Root Systems
Let $G$ be a connected reductive simple algebraic group over an algebraically closed field $K$ of simply connected type and rank $\ell$. Let $T$ be a maximal torus of $G$, $X\cong\mathbb{Z}^{\ell}$ its character group and $Y\cong\mathbb{Z}^{\ell}$ its co-character group. Let $\{\alpha_{1},\ldots,\alpha_{\ell}\}\subset X$ be a set of simple roots for $G$ with respect to $T$ and $\Delta$ its associated root system. Let $\alpha_{i}^{\vee}\in Y$ be the coroot corresponding to $\alpha_{i}$ for each $1\leq i\leq\ell$. Viewing $X\otimes\mathbb{R}$ as a Euclidean space we can define the fundamental weights $\lambda_{1}\ldots,\lambda_{\ell}\in X\otimes\mathbb{R}$ as the dual basis of $\alpha_{1}^{\vee},\ldots,\alpha_{\ell}^{\vee}$.
For our purposes, and in accordance with [24], we shall use the opposite order from that given in [6, 18]. Thus, the Dynkin diagrams for classical groups are given as follows:
[…]
We also recall that each $\alpha\in\Delta$ determines a reflection $\sigma_{\alpha}$ respect to the hyperplane $P_{\alpha}=\{\beta\in E:(\beta,\alpha)=0\}$, where (E,(,))(E,(-,-)) is the underlying inner product real vector space of finite dimensional associated to $\Delta$. The Weyl group $\mathcal{W}$ of $\Delta$ is the subgroup of the orthogonal group on $E$ generated by the reflections $\sigma_{\alpha_{i}}$. Thus, one observes that this group acts on $\Delta$ by permutation. In fact, $\mathcal{W}$ is generated by the reflections $\sigma_{\alpha_{i}}$ with $1\leq i\leq\ell$ by appealing to [20, Prop. 2.62].
Furthermore, one obtains a partial ordering in $X$ by declaring that a weight $\lambda$ is smaller than a weight $\lambda^{\prime}$ (denoted by $\lambda\prec\lambda^{\prime}$) if and only if $\lambda^{\prime}-\lambda$ is a non-negative linear combination of simple roots. We shall also use $\preceq$ in an obvious way. In addition, a weight $\lambda\in X$ is dominant if it is a non-negative linear combination of the fundamental weights. The group $\mathcal{W}$ acts on the set of the dominant weights by means of its action on the simple roots $\alpha_{1},\cdots,\alpha_{\ell}$. Given a dominant weight $\lambda$, we denote by $\operatorname{Orb}_{\mathcal{W}}(\lambda)$ and $\rm{Stab}_{\mathcal{W}}(\lambda)$ the orbit and the stabilizer of $\lambda$ under the $\mathcal{W}$-action, respectively.
It is worth recalling the following.
Remark 2.1.
The stabilizer subgroup $\rm{Stab}_{\mathcal{W}}(\lambda)$ of a dominant weight $\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}$ is the parabolic subgroup of $G$ generated by the reflections along the simple roots $\alpha_{i}$ such that $a_{i}=0$.
It will be used implicitly many times along the next sections mentioning it only the first instances we use it.
@@ 2.2. Representations
Henceforth, we assume that $K=\overline{\mathbb{F}}_{p}$ and let $V$ be a finite-dimensional $G(K)$-module. A classical result tells us
(2.1) $V=\bigoplus_{\mu\in X}V_{\mu},$
where $V_{\mu}:=\{v\in V:vt=\mu(t)v\mbox{ for all }t\in T\}$, called weight space, for any dominant weight $\mu\in X$. Each $\mu\in X$ which satisfies $V_{\mu}\neq\{0\}$ is called a weight of $V$. From a result of Chevalley (see Section 31.3 of [18]) we have that, when $V$ is irreducible, there exists a unique $T$-stable 11-dimensional subspace spanned by a unique maximal vector of dominant weight $\lambda$, is called the highest weight of $V$. Moreover, we have that each irreducible $G(K)$-module $V$ is determined (up to isomorphism) by its highest weight and for each dominant weight $\lambda\in X$ there exists an irreducible $G(K)$-module $L(\lambda)$ with highest weight $\lambda$. A way to construct $L(\lambda)$ is to start with a simple module $V(\lambda)_{\mathbb{C}}$ over $\mathbb{C}$ for the corresponding semisimple Lie algebra $\mathfrak{g}$. The theory in characteristic zero provides the existence of a special $\mathbb{Z}$-form $V(\lambda)_{\mathbb{Z}}\subset V(\lambda)_{\mathbb{C}}$, which allows reduction modulo $p$, such that $G$ acts naturally on the so-called Weyl module $V(\lambda):=K\otimes V(\lambda)_{\mathbb{Z}}$. Thus, the irreducible $G$-module $L(\lambda)$ is given by
$L(\lambda):=V(\lambda)/\rm{rad}(\lambda),$
where $\rm{rad}(\lambda)$ is the unique irreducible maximal submodule of $V(\lambda)$.
Let $F$ be the canonical Frobenius automorphism of $K$ given by $x\mapsto x^{p}$. Then, given a $G(K)$-module $L$ and $i\geq 0$, we denote by $L^{(i)}$ the $G(K)$-module obtained by composing the $G$-action on $L$ with the $i$-th power of $F$. On the other hand, we say that a dominant weight $\lambda=a_{1}\lambda_{1}+\cdots+a_{\ell}\lambda_{\ell}\in X$ is $p$-restricted if for every $i=1,\ldots,\ell$ we have $0\leq a_{i}\leq p-1$, and we may introduce the Twisted Tensor Product Theorem of Steinberg [33].
Theorem 2.2.
Let $\lambda\in X$ be a dominant weight. Then, there exist $r\in\mathbb{N}$ and $p$-restricted dominant weights $\lambda_{0},\cdots,\lambda_{r}\in X$ such that as $G(K)$-modules,
$L(\lambda)=L(\lambda_{0}+p\lambda_{1}+\cdots+p^{r}\lambda_{r})\cong L(\lambda_{0})\otimes L(\lambda_{1})^{(1)}\otimes\cdots\otimes L(\lambda_{r})^{(r)}.$
Then, the study of all irreducible $G(K)$-modules is reduced to that of the finitely many ones having $p$-restricted highest weights. In this vein, computing the dimension $L(\lambda)$ has been a theme of profound importance and multiple articles have contributed to obtain a better understanding of that. Next, we give a brief account of this matter.
2.3. Dimension and Multiplicity
We shall continue by presenting the well-known Freudenthal formula in characteristic zero. To this end, let $V_{\mathbb{C}}$ be a finite-dimensional irreducible $G$-module with highest weight $\lambda$. We need to introduce a special weight which appears in the computation of multiplicities as we will see here below. Specifically, the strongly dominant weight $\delta$ is defined by